Ben rolls a 1-6 number cube. Find each probability.
1.) Ben rolls a prime number or an odd number.
Mr. Rodney's English class is made up of 28 students. He has 6 ESL students make up 1/5 of the remedial students and 3/5 of the advanced learners.
2.)What is the probability that a student is remedial and NOT ESL?
Please explain how do you know.
Please answer me as soon as you could.
Thank you for your help.

Answer by stanbon(29602)

Ben rolls a 1-6 number cube. Find each probability.
1.) Ben rolls a prime number or an odd number.
---
P(prime or odd) = P(prime) + P(odd) - P(prime AND odd)
P(prime or odd) = 3/6 + 3/6 -- 2/6 = 4/6 = 2/3
------------------------------------------------------
Mr. Rodney's English class is made up of 28 students. He has 6 ESL students make up 1/5 of the remedial students and 3/5 of the advanced learners.
2.)What is the probability that a student is remedial and NOT ESL?
---
Your data is contradictory. If 6 ESL students make up 1/5 of the remedial
students there must be 30 remedial students in the class. But you say
there are a total of 28 students in the class. That information is
contradictory.
---------------------
Cheers,
Stan H.